### v1

@@ Abstract
We continue to study the learning-theoretic foundations of generation by extending the results from Kleinberg and Mullainathan (2024) and Li et al. (2024) to account for noisy example streams. In the noiseless setting of Kleinberg and Mullainathan (2024) and Li et al. (2024), an adversary picks a hypothesis from a binary hypothesis class and provides a generator with a sequence of its positive examples. The goal of the generator is to eventually output new, unseen positive examples. In the noisy setting, an adversary still picks a hypothesis and a sequence of its positive examples. But, before presenting the stream to the generator, the adversary inserts a finite number of negative examples. Unaware of which examples are noisy, the goal of the generator is to still eventually output new, unseen positive examples. In this paper, we provide necessary and sufficient conditions for when a binary hypothesis class can be noisily generatable. We provide such conditions with respect to various constraints on the number of distinct examples that need to be seen before perfect generation of positive examples. Interestingly, for finite and countable classes we show that generatability is largely unaffected by the presence of a finite number of noisy examples.
@@ 1 Introduction
Generation is an important paradigm in machine learning with promising applications to natural language processing (Wolf et al., 2020), computer vision (Khan et al., 2022), and computational chemistry (Vanhaelen et al., 2020). In contrast to its practical applications, a strong theoretical foundation of generation is largely missing from literature. Recently, Kleinberg and Mullainathan (2024), inspired by the seminal work of E Mark Gold on language identification in the limit (Gold, 1967), introduced a theoretical model of language generation called “Language Generation in the Limit.” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subset U$ for all $i\in\mathbbm{N}$. An adversary picks a language $K\in C$, and begins to enumerate the strings one by the one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses a string $\hat{w}_{t}\in U$ with the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. The class $C$ is generatable in the limit, if the player can generate from all $K\in C.$ Unlike language identification in the limit, Kleinberg and Mullainathan (2024) prove that generation in the limit is possible for every countable language family $C$.
@@
Kleinberg and Mullainathan (2024)’s positive result has spawned a surge of new work, extending the results of Kleinberg and Mullainathan (2024) in various ways (Kalavasis et al., 2024b; Li et al., 2024; Charikar and Pabbaraju, 2024; Kalavasis et al., 2024a). Of particular interest to us is the work by Li et al. (2024). They frame the results of Kleinberg and Mullainathan (2024) through a learning-theoretic lens by taking the language family $C$ to be a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over some countable instance space $\mathcal{X}.$ By doing so, Li et al. (2024) identify stronger notions of generatability, termed “uniform”  [^1] Uniform generatability was also informally considered by Kleinberg and Mullainathan (2024) and “non-uniform” generatability, and provide characterizations of which classes are uniformly and non-uniformly generatable in terms of new combinatorial dimensions.
@@
A commonality of Kleinberg and Mullainathan (2024) and its follow-up work is the assumption that the adversary presents to the player a “noiseless” stream of examples – one where every example/string must be contained in the hypothesis/language chosen by the adversary. In practice, such an assumption is unrealistic, as one would still like to generate well even if a few examples in the dataset are imperfect. For example, in light of recent interest in adversarial robustness, one might want a generative model to be robust to data contamination/poisoning attacks (Shang et al., 2018; Zhang et al., 2023; Jiang et al., 2023).
@@
Motivated by these concerns, we study generation, in the framework posed by Kleinberg and Mullainathan (2024) and Li et al. (2024), under noisy example streams. In particular, we focus on a simple, but natural noising process: after selecting a positive example stream, the adversary is allowed to insert a finite number of positive examples in any way it likes, unbeknownst to the player. Such a noising process, as well as others, have been extensively studied in the context “language identification in the limit” or “inductive inference” (Baliga et al., 1992; Fulk and Jain, 1989; Jain, 1994; Schäfer, 1985; Stephan, 1997; Mukouchi and Sato, 2003; Tantini et al., 2006; Lange and Grieser, 2002; Case et al., 1997). In this paper, we extend this study to generation. To that end, our main contributions are summarized below.
@@
    We extend the notions of uniform generatability, non-uniform generatability, and generatability in the limit from Kleinberg and Mullainathan (2024) and Li et al. (2024) to account for finite, noisy example streams. We call these new settings “noisy uniform generatability,” “noisy non-uniform generatability,” and “noisy generatability in the limit” respectively.
@@
    We provide a complete characterization of which classes are noisily uniformly generatable in terms of a new scale-sensitive dimension we call the Noisy Closure dimension.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily uniformly generatable if and only if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for every $n\in\mathbb{N}$, where $\operatorname{NC}_{n}(\mathcal{H})$ is the Noisy Closure dimension of $\mathcal{H}$ at noise-level $n.$
    We show that noisy uniform generation is strictly harder than noiseless uniform generation. In fact, we construct an countably infinite class which is trivially uniformly generatable when there is no noise, but not noisily uniformly generatable even when the adversary is only allowed to perturb a single example. Despite this hardness, we show that all finite classes are still noisily uniformly generatable.
@@
    We provide a sufficient condition for which classes are noisily non-uniformly generatable.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily non-uniformly generatable if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    Using the result that all finite classes are noisily uniformly generatable, we prove as a corollary that all countable classes are noisily non-uniformly generatable. Since noisy non-uniform generation implies noisy generation in the limit, this corollary also implies that all countable classes are noisily generatable in the limit. Although not matching, we also provide a necessary condition for noisy non-uniform generation in terms of the Noisy Closure dimension.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is noisily non-uniformly generatable only if for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    We leave the complete characterization of noisy non-uniform generatability as an open question.
@@
    Since Li et al. (2024) prove that all countable classes are (noiseless) non-uniform generatable, this result also shows that all countable classes are noisily generatable in the limit. In addition, we also give a sufficiency condition in terms in terms of a stronger notion of noisy generatability called “uniform noisy generatability.”
@@
    If there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
@@ 1.1 Related Work
Language Identification in the Limit. In his seminal 1967 paper, E Mark Gold introduced the model of “Language Identification in the Limit.” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subset U$ for all $i\in\mathbbm{N}$. An adversary picks a language $K\in C$, and begins to enumerate the strings in $K$ one by the one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses an index $i_{t}\in\mathbb{N}$ with the hope that $L_{i_{t}}=K$. The player has identified $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $L_{i_{s}}=K$. The class $C$ is identifiable in the limit, if the player can identify all $K\in C.$ Gold showed that while all finite language families are identifiable in the limit, there are simple countable language families which are not. In a follow-up work, Angluin (1979, 1980) provides a precise characterization of language families $C$ for which language identification in the limit possible. The results by Gold and Angluin emphasized the impossibility of language identification in the limit by ruling out the vast majority of language families. Since Gold’s seminal work on language identification in the limit, there has been extensive follow-up work on this model and we refer the reader to the excellent survey by Lange et al. (2008).
@@
Language Identification from Noisy Examples. In addition to the work by Angluin (1979, 1980), a different line of follow-up work focused on extending Gold’s model to account for noisy example streams. There have been several proposed noise models, and we review a few of them here. The most standard noise model for language identification in the limit allows the adversary to first pick an enumeration of its chosen language $K\in L$ and then insert a finite number of strings that do not belong to $K$ (Baliga et al., 1992; Fulk and Jain, 1989; Schäfer, 1985; Case et al., 1997; Stephan, 1997; Lange and Grieser, 2002). We use the term insert to emphasize that every $w\in K$ must still be present in the sequence chosen by the adversary. Such noisy streams are often referred to as noisy texts. Jain (1994) go beyond the finite nature of the noise by considering sequences of strings with an infinite number of noisy incorrect strings, where the amount of noise is measured using certain density notions. In a different direction, Mukouchi and Sato (2003) and Tantini et al. (2006) model noise at the string-level by defining a distance metric on the space of all strings and allowing the adversaries to insert and delete strings in the original enumeration of $K$ which are at most some distance away from some true string in $K$.
@@
Language Generation in the Limit. Inspired by large language models and recent interest in generative machine learning, Kleinberg and Mullainathan (2024) study the problem of language generation in the limit. In this problem, the adversary also picks a language $K\in C$, and begins to enumerate the strings one by the one to the player in rounds $t=1,2,\dots$. However, now, after observing the string $w_{t}$ in round $t\in\mathbbm{N}$, the player guesses a string $\hat{w}_{t}\in U$ with the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbbm{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. Kleinberg and Mullainathan (2024) prove a strikingly different result than Gold and Angluin – generation in the limit is possible for every countable language family $C$. This positive result has spurred a number of follow-up works, which we briefly review below.
@@
Kalavasis et al. (2024b) study generation in the stochastic setting, where the positive examples revealed to the generator are sampled i.i.d. from some unknown distribution. In this model, they study the trade-offs between generating with breadth and generating with consistency and show that in general, achieving both is impossible, resolving an open question posed by Kleinberg and Mullainathan (2024) for a large family of language models. More recently, Charikar and Pabbaraju (2024) and Kalavasis et al. (2024a) further formalize this tension between consistency and breadth by defining various notions of breadth and providing complete characterizations of which language families are generatable in the limit with breadth.
@@
In a different direction, and most closely related to our work, Li et al. (2024) reinterpret the results of Kleinberg and Mullainathan (2024) through a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over a countable example space $\mathcal{X}$. By doing so, Li et al. (2024) extend the results of Kleinberg and Mullainathan (2024) beyond language generation while also formalizing two stronger settings of generation they term “uniform” and “non-uniform” generation. Unlike Kleinberg and Mullainathan (2024) and Kalavasis et al. (2024a) who focus on computable learners and countable language families, Li et al. (2024) place no such restrictions and provide complete, information-theoretic characterizations of which hypothesis classes are uniformly and non-uniformly generatable.

### v2

@@ Abstract
We continue to study the learning-theoretic foundations of generation by extending the results from Kleinberg & Mullainathan (2024) and Li et al. (2024) to account for noisy example streams. In the noiseless setting of Kleinberg & Mullainathan (2024) and Li et al. (2024), an adversary picks a hypothesis from a binary hypothesis class and provides a generator with a sequence of its positive examples. The goal of the generator is to eventually output new, unseen positive examples. In the noisy setting, an adversary still picks a hypothesis and a sequence of its positive examples. But, before presenting the stream to the generator, the adversary inserts a finite number of negative examples. Unaware of which examples are noisy, the goal of the generator is to still eventually output new, unseen positive examples. In this paper, we provide necessary and sufficient conditions for when a binary hypothesis class can be noisily generatable. We provide such conditions with respect to various constraints on the number of distinct examples that need to be seen before perfect generation of positive examples. Interestingly, for finite and countable classes we show that generatability is largely unaffected by the presence of a finite number of noisy examples.
Machine Learning, ICML
@@ 1 Introduction
Generation is an important paradigm in machine learning with promising applications to natural language processing (Wolf et al., 2020), computer vision (Khan et al., 2022), and computational chemistry (Vanhaelen et al., 2020). In contrast to its practical applications, a strong theoretical foundation of generation is largely missing from literature. Recently, Kleinberg & Mullainathan (2024), inspired by the seminal work of E Mark Gold on language identification in the limit (Gold, 1967), introduced a theoretical model of language generation called “Language Generation in the Limit.” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subseteq U$ for all $i\in\mathbb{N}$. An adversary picks a language $K\in C$ and begins to enumerate the strings one by one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbb{N}$, the player guesses a string $\hat{w}_{t}\in U$ in the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbb{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. The class $C$ is generatable in the limit, if the player can generate from all $K\in C.$ Unlike language identification in the limit, Kleinberg & Mullainathan (2024) prove that generation in the limit is possible for every countable language family $C$.
@@
Kleinberg & Mullainathan (2024)’s positive result has spawned a surge of new work, extending the results of Kleinberg & Mullainathan (2024) in various ways (Kalavasis et al., 2024b; Li et al., 2024; Charikar & Pabbaraju, 2024; Kalavasis et al., 2024a). Of particular interest to us is the work by Li et al. (2024). They frame the results of Kleinberg & Mullainathan (2024) through a learning-theoretic lens by taking the language family $C$ to be a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over some countable instance space $\mathcal{X}.$ By doing so, Li et al. (2024) identify stronger notions of generatability, termed “uniform”  [^1] Uniform generatability was also informally considered by Kleinberg & Mullainathan (2024) and “non-uniform” generatability, and provide characterizations of which classes are uniformly and non-uniformly generatable in terms of new combinatorial dimensions.
@@
A commonality of Kleinberg & Mullainathan (2024) and its follow-up work is the assumption that the adversary presents to the player a “noiseless” stream of examples – one where every example/string must be contained in the hypothesis/language chosen by the adversary. In practice, such an assumption is unrealistic, as one would still like to generate well even if a few examples in the dataset are imperfect. For example, Large Language Models (LLM) are often trained using the potentially hallucinatory outputs of other LLMs (Burns et al., 2023; Briesch et al., 2023). More generally, one might want a generative model to be robust to data contamination/poisoning attacks (Shang et al., 2018; Zhang et al., 2023; Jiang et al., 2023).
@@
Motivated by these concerns, we study generation, in the framework posed by Kleinberg & Mullainathan (2024) and Li et al. (2024), under noisy example streams. In particular, we focus on a simple, but natural noising process: after selecting a positive example stream, the adversary is allowed to insert a finite number of negative examples in any way it likes, unbeknownst to the player. Such a noising process, as well as others, have been extensively studied in the context “language identification in the limit” or “inductive inference” (Schäfer, 1985; Fulk & Jain, 1989; Baliga et al., 1992; Jain, 1994; Stephan, 1997; Case et al., 1997; Lange & Grieser, 2002; Mukouchi & Sato, 2003; Tantini et al., 2006). In this paper, we extend this study to generation. To that end, our main contributions are summarized below.
@@
    We extend the notions of uniform generatability, non-uniform generatability, and generatability in the limit from Kleinberg & Mullainathan (2024) and Li et al. (2024) to account for finite, noisy example streams. We call these new settings “uniform noise-dependent generatability,” “non-uniform noise-dependent generatability,” and “noisy generatability in the limit” respectively.
@@
    We provide a complete characterization of which classes are uniformly noise-dependent generatable in terms of a new scale-sensitive dimension we call the Noisy Closure dimension.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is uniformly noise-dependent generatable if and only if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for every $n\in\mathbb{N}$, where $\operatorname{NC}_{n}(\mathcal{H})$ is the Noisy Closure dimension of $\mathcal{H}$ at noise-level $n.$
    We show that uniform noise-dependent generation is strictly harder than noiseless uniform generation. In fact, we construct a countably infinite class which is trivially uniformly generatable when there is no noise, but not uniformly noise-dependent generatable even when the adversary is only allowed to perturb a single example. Despite this hardness, we show that all finite classes are still uniformly noise-dependent generatable.
@@
    We provide a sufficient condition for which classes are non-uniformly noise-dependent generatable.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is non-uniformly noise-dependent generatable if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    Using the result that all finite classes are uniformly noise-dependent generatable, we prove as a corollary that all countable classes are non-uniformly noise-dependent generatable. Since non-uniform noise-dependent generation implies noisy generation in the limit, this corollary also implies that all countable classes are noisily generatable in the limit. Although not matching, we also provide a necessary condition for non-uniform noise-dependent generation in terms of the Noisy Closure dimension.
@@
    A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is non-uniformly noise-dependent generatable only if for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
    We leave the complete characterization of non-uniform noise-dependent generatability as an open question.
@@
    Since Li et al. (2024) prove that all countable classes are (noiseless) non-uniform generatable, this result also shows that all countable classes are noisily generatable in the limit. In addition, we also give a sufficiency condition in terms of an even stronger notion of noisy generatability called “uniform noise-independent generatability.”
@@
    If there exists a finite sequence of uniformly noise-independent generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
@@ 1.1 Related Work
Language Identification in the Limit. In his seminal 1967 paper, E Mark Gold introduced the model of “Language Identification in the Limit (Gold, 1967).” In this model, there is a countable set $U$ of strings and a countable language family $C=\{L_{1},L_{2},\dots\}$, where $L_{i}\subseteq U$ for all $i\in\mathbb{N}$. An adversary picks a language $K\in C$, and begins to enumerate the strings in $K$ one by the one to the player in rounds $t=1,2,\dots$. After observing the string $w_{t}$ in round $t\in\mathbb{N}$, the player guesses an index $i_{t}\in\mathbb{N}$ with the hope that $L_{i_{t}}=K$. The player has identified $K$ in the limit, if there exists a finite time step $t\in\mathbb{N}$ such that for all $s\geq t$, we have that $L_{i_{s}}=K$. The class $C$ is identifiable in the limit, if the player can identify all $K\in C.$ Gold showed that while all finite language families are identifiable in the limit, there are simple countable language families which are not. In a follow-up work, Angluin (1979, 1980) provides a precise characterization of language families $C$ for which language identification in the limit is possible. The results by Gold and Angluin emphasized the impossibility of language identification in the limit by ruling out the vast majority of language families. Since Gold’s seminal work on language identification in the limit, there has been extensive follow-up work on this model and we refer the reader to the excellent survey by Lange et al. (2008).
@@
Language Identification from Noisy Examples. In addition to the work by Angluin (1979, 1980), a different line of follow-up work focused on extending Gold’s model to account for noisy example streams. There have been several proposed noise models, and we review a few of them here. The most standard noise model for language identification in the limit allows the adversary to first pick an enumeration of its chosen language $K\in L$ and then insert a finite number of strings that do not belong to $K$ (Baliga et al., 1992; Fulk & Jain, 1989; Schäfer, 1985; Case et al., 1997; Stephan, 1997; Lange & Grieser, 2002). We use the term insert to emphasize that every $w\in K$ must still be present in the sequence chosen by the adversary. Such noisy streams are often referred to as noisy texts. Jain (1994) go beyond the finite nature of the noise by considering sequences of strings with an infinite number of noisy incorrect strings, where the amount of noise is measured using certain density notions. In a different direction, Mukouchi & Sato (2003) and Tantini et al. (2006) model noise at the string-level by defining a distance metric on the space of all strings and allowing the adversaries to insert and delete strings in the original enumeration of $K$ which are at most some distance away from some true string in $K$.
@@
Language Generation in the Limit. Inspired by large language models and recent interest in generative machine learning, Kleinberg & Mullainathan (2024) study the problem of language generation in the limit. In this problem, the adversary also picks a language $K\in C$, and begins to enumerate the strings one by the one to the player in rounds $t=1,2,\dots$. However, now, after observing the string $w_{t}$ in round $t\in\mathbb{N}$, the player guesses a string $\hat{w}_{t}\in U$ with the hope that $\hat{w}_{t}\in K\setminus\{w_{1},\dots,w_{t}\}$. The player has generated from $K$ in the limit, if there exists a finite time step $t\in\mathbb{N}$ such that for all $s\geq t$, we have that $\hat{w}_{s}\in K\setminus\{w_{1},\dots,w_{s}\}$. Kleinberg & Mullainathan (2024) prove a strikingly different result than Gold and Angluin – generation in the limit is possible for every countable language family $C$. This positive result has spurred a number of follow-up works, which we briefly review below.
@@
Kalavasis et al. (2024b) study generation in the stochastic setting, where the positive examples revealed to the generator are sampled i.i.d. from some unknown distribution. In this model, they study the trade-offs between generating with breadth and generating with consistency and show that in general, achieving both is impossible, resolving an open question posed by Kleinberg & Mullainathan (2024) for a large family of language models. More recently, Charikar & Pabbaraju (2024) and Kalavasis et al. (2024a) further formalize this tension between consistency and breadth by defining various notions of breadth and providing complete characterizations of which language families are generatable in the limit with breadth.
@@
In a different direction, and most closely related to our work, Li et al. (2024) reinterpret the results of Kleinberg & Mullainathan (2024) through a binary hypothesis class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ defined over a countable example space $\mathcal{X}$. By doing so, Li et al. (2024) extend the results of Kleinberg & Mullainathan (2024) beyond language generation while also formalizing two stronger settings of generation they term “uniform” and “non-uniform” generation. Unlike Kleinberg & Mullainathan (2024) and Kalavasis et al. (2024a) who focus on computable learners and countable language families, Li et al. (2024) place no such restrictions and provide complete, information-theoretic characterizations of which hypothesis classes are uniformly and non-uniformly generatable. Li et al. (2024) leave the characterization of generatability in the limit as an open question (see Section 4).